OpenStudy (anonymous):

Part 1: Solve the inequality 2(x + 6) >= 3x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

2x+12 >= 3x-2 14 >= x x <= 14 The graph will have a closed circle on x = 14 and will shade left towards -infty

OpenStudy (anonymous):

-infinity*

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

2(x+6)>=3x-2 2x+12>=3x-2 3x-2x<=12+2 x<=14 close circle on 14:shading to the left